# From Image Denoisers to Regularizing Imaging Inverse Problems: An Overview

### v1

@@ 2.3 Accelerated Convex Methods
From the previous reformulation of plug-and-play methods as non-convex optimization of some explicit functionals, a natural question is whether or not the optimization can be accelerated. This has direct implications for the reconstruction speed, which is given by the rate of convergence to critical points of said functionals. There are two main ways of acceleration in the optimization literature, namely via momentum and preconditioning. We note that there are no available convergence guarantees in the former case of momentum-based accelerated PnP, other than spectral analyses for linear denoisers and linear inverse problems sinha2025fista . In gradient-based optimization, preconditioning refers to multiplying the gradient by some (usually positive definite) preconditioner matrix, to make the problem “less ill-conditioned” and achieve faster convergence using larger step sizes. Common instances include Newton’s method, Riemannian gradient descent, or the more exotic mirror descent. In the proximal splitting case, however, the preconditioning affects not only the gradient step but also the proximal step. For example, the preconditioned proximal gradient method to minimize $f+g$ takes the form
$x_{k+1}=\operatorname{prox}_{g}^{B_{k}}(x-B_{k}^{-1}\nabla f(x_{k})),$ (25a)
@@
In hong2024provable , the authors consider both static and dynamic preconditioners, with applications to MRI reconstruction. For the dynamic preconditioning, convergence of the variable proximal scaling assumes a “normalization-equivariant denoiser” $D$, which is assumed to have the property that for any $\mu>0$ and $\Delta\in\mathbb{C}$, that $D(\mu x+\Delta\mathbf{1})=\mu D(x)+\Delta\mathbf{1}$. The idea of adding an adjustable denoiser with diagonal preconditioners to PnP-ADMM is considered in pendu2023preconditioned , with applications to Poisson denoising.
@@ 3 Learning Image Priors using Denoisers
In this section, we will discuss how denoisers can be used to construct an explicit prior, in contrast with proximal PnP schemes where denoisers provide implicit regularization. The main theoretical foundation for designing direct regularizers from Gaussian image denoisers rests on Tweedie’s formula tweedie_main .
@@ 3.1 Tweedie’s Formula
The use of Gaussian image denoisers as priors can be justified through Tweedie’s formula. Consider the task of estimating the clean image $x$ with a probability density function (p.d.f.) $p(x)$ from its noisy measurement $x_{\sigma}=x+\sigma\,w$, where $\sigma>0$ is the noise standard deviation and $w\sim\mathcal{N}(0,\text{Id})$. The p.d.f. of the noisy image $x_{\sigma}$ is given by
@@
Rearranging the terms in (30) leads to the familiar Tweedie’s formula:
$\mathbb{E}\left[x|x_{\sigma}\right]-x_{\sigma}=\sigma^{2}\,\nabla_{x_{\sigma}}\log p_{\sigma}(x_{\sigma}),$ (31)
which establishes a direct connection between the optimal MMSE Gaussian denoiser and the score function $\nabla_{x_{\sigma}}\log p_{\sigma}(x_{\sigma})$ of the noisy image $x_{\sigma}$. The optimal Gaussian image denoiser essentially seeks to approximate the conditional mean of the clean image given its noisy observation (as given in (31)), and hence provides a way for approximating the noisy score function. Such a denoiser is learned by minimizing (an empirical approximation of) the MSE loss $\mathbb{E}_{x,x_{\sigma}}\left\|D(x_{\sigma})-x\right\|_{2}^{2}$ on a training dataset having clean and noisy image pairs.
@@ 3.2 Regularization-by-Denoising (RED)
$x_{k+1}=x_{k}-\eta[K^{*}(Kx_{k}-y)+\frac{\lambda}{\sigma^{2}}(x_{k}-D(x_{k}))].$ (33)
Compared to PnP-PGD, which takes the denoiser as an implicit proximal operator $x_{k+1}=D_{\sigma}(x_{k}-\eta\,K^{*}(Kx_{k}-y))$, using 33 allows for a more intuitive practical tuning for RED. As reported in romano2017little , proximal PnP’s performance is highly dependent on other algorithmic parameters such as the step size $\eta$ in PnP-PGD. However, for RED, as long as the regularization parameter $\lambda$ is selected, RED will converge to fixed points with very similar reconstruction quality, as long as the step size $\eta$ is below a certain threshold.
@@
If the denoiser $D$ has a asymmetric Jacobian, then there is no regularization $R(x)$ that satisfies $\nabla R(x)=x-D(x)$.
@@ 3.3 Convergence Theory for RED
$K^{*}(Kx^{*}-y)+\lambda\left(x^{*}-D(x^{*})\right)=0.$ (38)
To demonstrate, we consider a provably convergent RED algorithm, namely proximal gradient RED (RED-PG) reehorst2019regularization . Given a data fidelity $f(x)$ such as the least-squares error $\|Kx-y\|_{2}^{2}$, parameters $\lambda>0$, initialization $v_{0}$, $L>0$, the iterations of RED-PG are described as follows:
@@
Using the link between Tweedie’s formula and RED 32, the term inside (the argument of the proximal operator in (40)) may be interpreted as an approximate gradient ascent step on the log-prior, since
@@ 4 Convergence of Proximal PnP Methods
For stability and mathematical interpretation, it is important to guarantee the convergence of PnP iterations. Since (proximal) PnP methods are derived by replacing proximal operators using off-the-shelf denoisers, their convergence is not automatically guaranteed with a generic denoiser. In this section, we will define some key notions of convergence for PnP methods from weak to strong, and highlight their practical significance. We will also present some recent representative foundational convergence theorems under each category.
@@ 4.1 Fixed-Point/Iterate Convergence
This is by far the weakest form of convergence, which requires only that the PnP iterations converges to a solution. More formally, the studies on fixed-point convergence consider the PnP iterations as a fixed-point update rule of the form $x_{k+1}=\operatorname{\mathcal{T}}(x_{k})$, and seek to determine whether $x_{k}$ converges to a fixed-point $x^{*}$ of $\operatorname{\mathcal{T}}$. The specific structure of the operator $\operatorname{\mathcal{T}}$ depends on the choice of the proximal splitting algorithm, the forward operator, and the denoiser. For instance, for PnP-PGD, the iterations are given by $x_{k+1}=D\left(x_{k}-\eta\,\nabla f(x_{k})\right)$, resulting in $\operatorname{\mathcal{T}}=D\circ\left(\operatorname{Id}-\eta\nabla f\right)$. We say that a PnP method is fixed-point convergent if $\operatorname{\mathcal{T}}$ has a unique fixed-point $x^{*}$ (i.e., $\operatorname{\mathcal{T}}(x^{*})=x^{*}$) and the PnP iterations converge to $x^{*}$, meaning that $\underset{k\rightarrow\infty}{\lim}\,\,x_{k}=x^{*}$. Achieving fixed-point convergence of a PnP algorithm essentially boils down to ensuring that $\operatorname{\mathcal{T}}$ is a contraction mapping (under suitable conditions on the forward operator and the denoiser). This mode of convergence inherently guarantees that the solution does not worsen as the PnP iterations are repeated a large number of times.
Theorem 4.1 (Fixed-point convergence of PnP-DRSdiff (ryu2019plug, , Thm. 4.3))
@@
$x_{k+\frac{1}{2}}$ $=\operatorname{prox}_{\tau f}\left(z_{k}\right),\quad x_{k+1}=D\left(2x_{k+\frac{1}{2}}-z_{k}\right),\text{\,\,and\,\,}$
$z_{k+1}$ $=z_{k}+x_{k+1}-x_{k+\frac{1}{2}},$ (41)
where the data-fidelity term $f$ is $\mu$-strongly convex. One can equivalently express (41) as a fixed-point iteration of the form $z_{k+1}=\operatorname{\mathcal{T}}(z_{k})$, with
$\operatorname{\mathcal{T}}=\frac{1}{2}\operatorname{Id}+\frac{1}{2}\left(2D-\operatorname{Id}\right)\left(2\,\operatorname{prox}_{\tau f}-\operatorname{Id}\right).$ (42)
Suppose that the denoiser $D$ satisfies
@@
for all $u,v\in\mathbb{X}$ and some $\epsilon>0$, and the strong convexity parameter $\mu$ is such that $\frac{\epsilon}{(1+\epsilon-2\epsilon^{2})\,\mu}<\tau$ is satisfied. Then the operator $\operatorname{\mathcal{T}}$ is contractive and the PnP-ADMM algorithm is fixed-point convergent. That is, $\left(x_{k},z_{k}\right)\rightarrow(x^{*},z^{*})$, where $(x^{*},z^{*})$ satisfy
$x^{*}=\operatorname{prox}_{\tau f}\left(z^{*}\right)\text{\,\,and\,\,}x^{*}=D\left(2x^{*}-z^{*}\right).$ (44)
@@
4.2 Objective Convergence
While convergence to a fixed point guarantees stability under repeated iterations, the fixed point generally does not lend itself to a variational interpretation. That is, the fixed point is not generally a minimizer or a stationary point of a variational energy function induced by the denoiser. Objective convergence of PnP draws a direct parallel between PnP and classical variational schemes by ensuring that the PnP solution is indeed a stationary point of some variational objective (which can potentially be non-convex depending on the denoiser). Such convergence can be shown by imposing special structures on the denoiser (while leveraging convergence analysis of proximal gradient descent or some variant of it in the non-convex setting), as we state in the theorem below.
Theorem 4.3 (Objective convergence of PnP iterations with gradient-step (GS) denoisers gs_denoiser_hurault_2021 )
Suppose the denoiser is constructed as a gradient-step (GS) denoiser, i.e., $D=\operatorname{Id}-\nabla g$, where $g$ is proper, lower semi-continuous, and differentiable with an $L$-Lipschitz gradient. The gradient-step PnP algorithm proposed in gs_denoiser_hurault_2021 is given by
$=\operatorname{prox}_{\tau f}\circ\left(\tau\lambda\,D+(1-\tau\lambda\,\operatorname{Id})\right)(x_{k}),$ (45)
where $f\colon X\to\mathbb{R}\cup\{+\infty\}$ denotes data-fidelity and is assumed to be convex and lower semi-continuous. Then, the following guarantees hold for $\tau<\frac{1}{\lambda\,L}$:
  1. 1.
  2. 2.
  3. 3.
    All limit points of $\{x_{k}\}$ are stationary points of $F(x)$.
The construction of gradient-step denoisers is motivated by Tweedie’s formula, which states that the optimal minimum mean squared-error (MMSE) Gaussian denoiser indeed has the form $D(x)=\operatorname{Id}-\nabla g(x)$. Similarly to red_romano , $g(x)$ is constructed as $g(x)=\frac{1}{2}\left\|x-N(x)\right\|_{2}^{2}$, where $N(x)$ is taken to be the DRUNet zhang2021plug . As noted in gs_denoiser_hurault_2021 , this specific design leads to a powerful denoiser while facilitating convergence analysis. The GS-PnP algorithm and following analysis can be seen to be similar to PnP-PGD, with the gradient step and proximal step flipped. Some other recent PnP objective convergence results under specific technical assumptions on the denoiser can be found in kunal_9380942 ; tan2024provably .
4.3 Convergent Regularization using PnP
While objective convergence ensures a one-to-one connection between PnP iterates and the minimization of a variational objective, it does not provide any guarantees about the regularizing properties of the solution that the iterates converge to. In the same spirit as classical regularization theory, it is therefore desirable to be able to control the implicit regularization effected by the denoiser in PnP algorithms and analyze the limiting behavior of the PnP reconstruction as the noise level and the regularization strength tend to zero. More precisely, assuming that the PnP iterations converge to a solution $\hat{x}\left(y^{\delta},\lambda\right)$, where $\delta$ denotes the noise level and $\lambda$ is an explicit regularization penalty parameter associated with the denoiser, one would like to obtain appropriate selection rules for $\sigma$ and/or $\lambda$ such that $\hat{x}\left(y^{\delta},\lambda\right)$ exhibits convergence akin to (5) in the limit as $\delta\rightarrow 0$. To the best of our knowledge, one of the first analyses of this kind was reported in ebner2022plugandplay , and the precise convergence result is stated in Theorem 4.4.
Theorem 4.4 (Convergent PnP regularization ebner2022plugandplay )
$x_{\lambda,k+1}^{\delta}=D_{\lambda}\left(x_{\lambda,k}^{\delta}-\eta\,K^{*}\left(Kx_{\lambda,k}^{\delta}-y^{\delta}\right)\right),$ (46)
where $D_{\lambda}$ is a denoiser with a tuneable regularization parameter $\lambda$. Let $\operatorname{PnP}\left(\lambda,y^{\delta}\right)$ be the fixed point of the PnP iteration (46). For any $y\in\mathscr{R}(K)$ and any sequence $\delta_{k}>0$ of noise levels converging to 0, there exists a sequence $\lambda_{k}$ of regularization parameters converging to 0 such that for all $y_{k}$ with $\|y_{k}-y^{0}\|_{2}\leq\delta_{k}$, the following hold under appropriate assumptions on the denoiser (see Definition 3.1 in ebner2022plugandplay for details):
  1. 1.
    $\operatorname{PnP}\left(\lambda,y^{\delta}\right)$ is continuous in $y^{\delta}$ for any $\lambda>0$;
  2. 2.
    The sequence $\left(\operatorname{PnP}\left(\lambda_{k},y_{k}\right)\right)_{k\in\mathbb{N}}$ has a weakly convergent subsequence; and
  3. 3.
    The limit of every weakly convergent subsequence of $\left(\operatorname{PnP}\left(\lambda_{k},y_{k}\right)\right)_{k\in\mathbb{N}}$ is a solution of the noiseless operator equation $y^{0}=Ax$.
5 Practical Constraints and Training
5.1 Weakly Enforced Spectral Constraints
For convergence analysis, one key requirement is that the denoiser should take the form of a proximal step. For a gradient step denoiser $D_{\sigma}=I-\nabla g_{\sigma}$ to be a proximal operator $D_{\sigma}=\operatorname{prox}_{\phi_{\sigma}}$ of some weakly convex function $\phi_{\sigma}$, a sufficient condition is that $\nabla g_{\sigma}$ is $L_{\sigma}$-Lipschitz for some $L_{\sigma}<1$ gribonval2020characterization ; hurault2022proximal . Here, we use the subscript $\sigma$ to denote specifically that the denoiser is trained to remove Gaussian noise of standard deviation $\sigma$. As strongly enforcing the Lipschitz condition is difficult, a standard approach in practice is to consider using a soft penalty in the loss function. Noting the equivalence of the Lipschitz constant and the spectral norm of $\nabla^{2}g_{\sigma}=J(I-D_{\sigma})$, this consists of adding a spectral regularization term of the form
$\mathbb{E}_{x\sim p,\xi\sim\mathcal{N}(0,\sigma^{2})}\max(\|J(I-D_{\sigma})(x+\xi)\|,1-\varepsilon).$
This penalizes the spectral norm of $\nabla^{2}g_{\sigma}$, typically approximated using a power iteration. Other methods of softly enforcing the Lipschitz constant include (approximate) layer-wise projections onto the Stiefel manifold of orthogonal matrices anil2019sorting ; cisse2017parseval , or doing the same with monotone operators pesquet2021learning ; belkouchi2025learning . However, as the Lipschitz constant is not strictly enforced to be less than one, the algorithms suffer from occasional divergence.
5.2 Backtracking for Lipschitz Control
As mentioned in Theorem˜4.3, the gradient-step paradigm for PnP instead replaces the gradient step in a splitting with a denoiser, and applies the proximal operator on the fidelity term gs_denoiser_hurault_2021 . In this case, since the step size in the splitting is allowed to be variable, it remains to find an upper bound on the Lipschitz constant of $\nabla g_{\sigma}$, such that $D_{\lambda,\sigma}=I-\lambda\nabla g_{\sigma}$ is a descent step.
@@ 5.3 Kurdyka–Łojasiewicz Property
Another weaker form of convergence utilizes the Kurdyka–Łojasiewicz (KL) property. This is a general property that may be satisfied using architectural choices on neural networks. In the following, $\partial^{l}$ denotes the limiting subdifferential.
@@
The KL property can be interpreted as a certain regularity condition. Indeed, if $\varphi$ satisfies the KL property at a critical point $\overline{u}$, then it can be shown that subgradients of $u\mapsto\Psi\circ(\varphi(u)-\varphi(\overline{u}))$ have norm bounded away from one, also known as sharpness bolte2014proximal . This geometric property ensures that critical points have sufficient regularity properties. While the definition is seemingly complicated, the KL property holds for many classes of functions, with some examples given in the following proposition.
@@
The (sub)analytic functions and semialgebraic characterizations are particularly useful. They imply that smooth neural networks and networks with piecewise-polynomial activations, such as ReLU, both satisfy the KL property, which can be used to show convergence of PnP methods in nonconvex settings. In the absence of convexity, one can use the KL property to show convergence. For example, the following abstract theorem shows convergence under certain conditions on the iterates. Moreover, the convergence is fast in the sense that the sequence has finite length.
@@
If, furthermore, $f$ has the KL property at the cluster point $\tilde{x}$, then the entire sequence $(x_{k})_{k\in\mathbb{N}}$ converges to $\tilde{x}$. Moreover, $\tilde{x}$ is a critical point of $f$, and $(x_{k})_{k\in\mathbb{N}}$ has finite length, i.e., $\sum_{k\in\mathbb{N}}\|x_{k+1}-x_{k}\|<+\infty$.

### v2

@@ 2.3 Accelerated Convex Methods
From the previous reformulation of plug-and-play methods as non-convex optimization of some explicit functionals, a natural question is whether or not the optimization and therefore reconstruction can be accelerated. There are two main ways of acceleration in the optimization literature, namely via momentum and preconditioning. We note that there are no available convergence guarantees in the former case of momentum-based accelerated PnP, other than spectral analyses for linear denoisers and linear inverse problems sinha2025fista . In gradient-based optimization, preconditioning refers to multiplying the gradient by some (usually positive definite) preconditioner matrix, to make the problem “less ill-conditioned” and achieve faster convergence using larger step sizes. Common instances include Newton’s method, Riemannian gradient descent, or the more exotic mirror descent. In the proximal splitting case, however, the preconditioning affects not only the gradient step but also the proximal step. For example, the preconditioned proximal gradient method to minimize $f+g$ takes the form
$x_{k+1}=\operatorname{prox}_{g}^{B_{k}}(x_{k}-B_{k}^{-1}\nabla f(x_{k})),$ (25a)
@@
In hong2024provable , the authors consider both fixed and iteration-dependent preconditioners, with applications to MRI reconstruction. For the iteration-dependent preconditioning, convergence of the variable proximal scaling assumes a “normalization-equivariant denoiser” $D$, which is assumed to have the property that for any $\mu>0$ and $\Delta\in\mathbb{C}$, that $D(\mu x+\Delta\mathbf{1})=\mu D(x)+\Delta\mathbf{1}$ where $\mathbf{1}$ represents the vector of all ones. Denoisers can also be made to be “adjustable”, by allowing them to take an additional input corresponding to the preconditioner. An application with diagonal preconditioners to PnP-ADMM is considered in pendu2023preconditioned , with applications to Poisson denoising.
@@ 3 Learning Image Priors using Denoisers
In this section, we will discuss how denoisers can be used to construct an explicit prior, in contrast with proximal PnP schemes where denoisers provide implicit regularization. In the case where denoisers are used to target Gaussian noise, henceforth known as Gaussian image denoisers, a theoretical link towards Bayesian inference can be drawn using Tweedie’s formula tweedie_main .
@@ 3.1 Tweedie’s Formula
Consider the task of estimating the clean image $x$ with a probability density function (p.d.f.) $p(x)$ from its noisy measurement $x_{\sigma}=x+\sigma\,w$, where $\sigma>0$ is the noise standard deviation and $w\sim\mathcal{N}(0,\text{Id})$. The p.d.f. of the noisy image $x_{\sigma}$ is given by
@@
Rearranging the terms in (30) leads to the familiar Tweedie’s formula, a direct connection between the optimal minimum mean squared error (MMSE) Gaussian denoiser $\mathbb{E}[x|x_{\sigma}]$, and the score function $\nabla_{x_{\sigma}}\log p_{\sigma}(x_{\sigma})$ of the noisy image $x_{\sigma}$:
$\mathbb{E}\left[x|x_{\sigma}\right]-x_{\sigma}=\sigma^{2}\,\nabla_{x_{\sigma}}\log p_{\sigma}(x_{\sigma}).$ (31)
The optimal Gaussian image denoiser essentially seeks to approximate the conditional mean of the clean image given its noisy observation (as given in (31)), and hence provides a way for approximating the noisy score function. Such a denoiser is learned by minimizing (an empirical approximation of) the MSE loss $\mathbb{E}_{x,x_{\sigma}}\left\|D(x_{\sigma})-x\right\|_{2}^{2}$ on a training dataset having clean and noisy image pairs.
@@ 3.2 Regularization-by-Denoising (RED)
$x_{k+1}=x_{k}-\eta[K^{\top}(Kx_{k}-y)+\frac{\lambda}{\sigma^{2}}(x_{k}-D(x_{k}))].$ (33)
@@
If the denoiser $D$ has an asymmetric Jacobian, then there is no regularization $R(x)$ that satisfies $\nabla R(x)=x-D(x)$.
@@ 3.3 Convergence Theory for RED
$K^{\top}(Kx^{*}-y)+\lambda\left(x^{*}-D(x^{*})\right)=0.$ (38)
To demonstrate convergence, we consider a provably convergent RED algorithm, namely proximal gradient RED (RED-PG) reehorst2019regularization . Given a data fidelity $f(x)$ such as the least-squares error $\|Kx-y\|_{2}^{2}$, parameters $\lambda>0$, $L>0$, and initialization $v_{0}$, the iterations of RED-PG are described as follows:
@@
Using the link between Tweedie’s formula and RED 32, the argument of the proximal may be interpreted as an approximate gradient ascent step on the log-prior, since
@@ 4 Convergence of Proximal PnP Methods
For stability and mathematical interpretation, it is important to guarantee the convergence of PnP iterations. Since (proximal) PnP methods are derived by replacing proximal operators using off-the-shelf denoisers, their convergence is not automatically guaranteed with a generic denoiser. In this section, we will define some key notions of convergence for PnP methods from weak to strong, and highlight their practical significance. We will also present some recent representative foundational convergence theorems under each category; for additional results under each category, we refer to Table˜1 and towards the references.
@@ 4.1 Fixed-Point/Iterate Convergence
This is by far the weakest form of convergence, which requires only that the PnP iterations converge to a solution of some fixed point problem. More formally, the studies on fixed-point convergence consider the PnP iterations as a fixed-point update rule of the form $x_{k+1}=\operatorname{\mathcal{T}}(x_{k})$, and seek to determine whether $x_{k}$ converges to a fixed-point $x^{*}$ of $\operatorname{\mathcal{T}}$. The specific structure of the operator $\operatorname{\mathcal{T}}$ depends on the choice of the proximal splitting algorithm, the forward operator, and the denoiser. For instance, for PnP-PGD, the iterations are given by $x_{k+1}=D\left(x_{k}-\eta\,\nabla f(x_{k})\right)$, resulting in $\operatorname{\mathcal{T}}=D\circ\left(\operatorname{Id}-\eta\nabla f\right)$. We say that a PnP method is fixed-point convergent if $\operatorname{\mathcal{T}}$ has a unique fixed-point $x^{*}$ (i.e., $\operatorname{\mathcal{T}}(x^{*})=x^{*}$) and the PnP iterations converge to $x^{*}$, meaning that $\underset{k\rightarrow\infty}{\lim}\,\,x_{k}=x^{*}$. Achieving fixed-point convergence of a PnP algorithm essentially boils down to ensuring that $\operatorname{\mathcal{T}}$ is a contraction mapping (under suitable conditions on the forward operator and the denoiser). This mode of convergence inherently guarantees that the solution does not worsen as the PnP iterations are repeated a large number of times.
Theorem 4.1 (Fixed-point convergence of PnP-DRSdiff (ryu2019plug, , Thm. 3))
@@
$\begin{cases}y_{k}=\operatorname{prox}_{\tau f}\left(x_{k}\right),\\ z_{k}=D(2y_{k}-x_{k}),\\ x_{k+1}=x_{k}+z_{k}-y_{k},\end{cases}$ (41)
where the data-fidelity term $f$ is $\mu$-strongly convex. Letting $\operatorname{\mathcal{T}}$ be the following operator
$\operatorname{\mathcal{T}}=\frac{1}{2}\operatorname{Id}+\frac{1}{2}\left(2D-\operatorname{Id}\right)\left(2\operatorname{prox}_{\tau f}-\operatorname{Id}\right),$ (42)
one may equivalently express (41) as a fixed-point iteration of the form $x_{k+1}=\operatorname{\mathcal{T}}(x_{k})$. Suppose that the denoiser $D$ satisfies
@@
for all $u,v\in\mathcal{X}$ and some $\epsilon>0$, and the strong convexity parameter $\mu$ is such that $\frac{\epsilon}{(1+\epsilon-2\epsilon^{2})\,\mu}<\tau$ is satisfied. Then the operator $\operatorname{\mathcal{T}}$ is contractive and the PnP-DRSdiff algorithm is fixed-point convergent.
@@ 4.2 Kurdyka–Łojasiewicz Property
In the absence of convexity of the prior or contractivity of the denoiser, another weaker form of convergence utilizes the Kurdyka–Łojasiewicz (KL) property of a function. This is a general property that may be satisfied using architectural choices on neural networks. In the following, $\partial^{l}$ denotes the limiting subdifferential.
@@
The KL property can be interpreted as a certain regularity condition on a function $\varphi$. Indeed, if $\varphi$ satisfies the KL property at a critical point $\overline{u}$, then it can be shown that subgradients of $u\mapsto\Psi(\varphi(u)-\varphi(\overline{u}))$ have norm bounded away from one, also known as sharpness bolte2014proximal . This geometric property ensures that critical points have sufficient regularity properties. While the definition is seemingly complicated, the KL property holds for many classes of functions, with some examples given in the following proposition.
@@
The (sub)analytic functions and semialgebraic characterizations are particularly useful. As a special case, smooth neural networks and networks with piecewise-polynomial activations, such as ReLU, both satisfy the KL property. In the absence of convexity, one can instead use the KL property to show convergence. For example, the following abstract theorem shows convergence under certain conditions on the iterates. Moreover, the convergence is fast in the sense that the sequence has finite length.
@@
If, furthermore, $f$ has the KL property at the cluster point $\tilde{x}$, then the entire sequence $(x_{k})_{k\in\mathbb{N}}$ converges to $\tilde{x}$. Moreover, $\tilde{x}$ is a critical point of $f$, and $(x_{k})_{k\in\mathbb{N}}$ has finite length.
@@
4.3 Objective Convergence
While convergence to a fixed point guarantees stability under repeated iterations, the fixed point generally does not lend itself to a variational interpretation. That is, the fixed point is not generally a minimizer or a stationary point of a variational energy function induced by the denoiser. Objective convergence of PnP draws a direct parallel between PnP and classical variational schemes by ensuring that the PnP solution is indeed a stationary point of some variational objective (which can potentially be non-convex depending on the denoiser). Such convergence can be shown by imposing special structures on the denoiser (while leveraging convergence analysis of proximal gradient descent or some variant of it in the non-convex setting).
One popular structure is to define the denoiser as a gradient-step (GS), result in a so-called GS denoiser. These denoisers take the form $D=\operatorname{Id}-\nabla g$, where the “potential function” $g$ is proper, lower semi-continuous, and differentiable with an $L$-Lipschitz gradient. Using this structure, the denoiser can be substituted into the gradient step of the FBS: for a step-size $\tau>0$ and relaxation parameter $\lambda>0$,
$=\operatorname{prox}_{\tau f}\circ\left(\tau\lambda\,D+(1-\tau\lambda)\operatorname{Id}\right)(x_{k}),$ (44)
where $\circ$ denotes function composition. GS denoisers enjoy the following convergence.
Theorem 4.6 (Objective convergence of PnP iterations with GS denoisers gs_denoiser_hurault_2021 )
Suppose the denoiser is constructed as a GS denoiser $D=\operatorname{Id}-\nabla g$, where $g$ is proper, lower semi-continuous, and differentiable with an $L$-Lipschitz gradient. Suppose further that the data-fidelity $f\colon\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$ is convex and lower semi-continuous. Then, the following guarantees hold for $\tau<\frac{1}{\lambda\,L}$:
  1. 1.
  2. 2.
  3. 3.
    All limit points of $\{x_{k}\}$ are stationary points of $F(x)$.
The construction of gradient-step denoisers is motivated by Tweedie’s formula, which states that the optimal minimum mean squared-error (MMSE) Gaussian denoiser indeed has the form of a gradient step under some prior term $D(x)=\operatorname{Id}-\nabla g(x)$. Similarly to romano2017little , the potential $g(x)$ typically takes the form $g(x)=\frac{1}{2}\left\|x-N(x)\right\|_{2}^{2}$, where $N(x)$ is taken to be some differentiable neural network such as DRUNet zhang2021plug . As noted in gs_denoiser_hurault_2021 , this specific design leads to a powerful denoiser while facilitating convergence analysis. The GS-PnP algorithm and following analysis can be seen to be similar to PnP-PGD, with the gradient step and proximal step flipped. Some other recent PnP objective convergence results under specific technical assumptions on the denoiser can be found in kunal_9380942 ; tan2024provably .
4.4 Convergent Regularization Using PnP
While objective convergence ensures a one-to-one connection between PnP iterates and the minimization of a variational objective, it does not provide any guarantees about the regularizing properties of the solution that the iterates converge to. In the same spirit as classical regularization theory, it is therefore desirable to be able to control the implicit regularization induced by the denoiser in PnP algorithms, and analyze the limiting behavior of the PnP reconstruction as the noise level and the regularization strength tend to zero. More precisely, assuming that the PnP iterations converge to a solution $\hat{x}\left(y^{\delta},\lambda\right)$, where $\delta$ denotes the noise level and $\lambda$ is an explicit regularization penalty parameter associated with the denoiser, one would like to obtain appropriate selection rules for $\sigma$ and/or $\lambda$ such that $\hat{x}\left(y^{\delta},\lambda\right)$ exhibits convergence akin to (5) in the limit as $\delta\rightarrow 0$. To the best of our knowledge, one of the first analyses of this kind was reported in ebner2022plugandplay , and the precise convergence result is stated in Theorem 4.7.
Theorem 4.7 (Convergent PnP regularization (ebner2022plugandplay, , Thm. 3.14))
$x_{\lambda,k+1}^{\delta}=D_{\lambda}\left(x_{\lambda,k}^{\delta}-\eta\,K^{\top}\left(Kx_{\lambda,k}^{\delta}-y^{\delta}\right)\right),$ (45)
where $D_{\lambda}$ is a denoiser with a tuneable regularization parameter $\lambda$. Suppose that the family of denoisers $\{D_{\lambda}\}_{\lambda>0}$ satisfies appropriate assumptions (see Definition 3.1 in ebner2022plugandplay for details), in particular that they are contractive so that the PnP iterations converge. Let $\operatorname{PnP}\left(\lambda,y^{\delta}\right)$ be the fixed point of the PnP iteration (45). For any $y=y^{0}\in\mathscr{R}(K)$ and any sequence $\delta_{k}>0$ of noise levels converging to 0, there exists a sequence $\lambda_{k}$ of regularization parameters converging to 0 such that for all $y_{k}$ with $\|y_{k}-y^{0}\|_{2}\leq\delta_{k}$:
  1. 1.
    $\operatorname{PnP}\left(\lambda,y^{\delta}\right)$ is continuous in $y^{\delta}$ for any $\lambda>0$.
  2. 2.
    The sequence $\left(\operatorname{PnP}\left(\lambda_{k},y_{k}\right)\right)_{k\in\mathbb{N}}$ has a weakly convergent subsequence.
  3. 3.
    The limit of every weakly convergent subsequence of $\left(\operatorname{PnP}\left(\lambda_{k},y_{k}\right)\right)_{k\in\mathbb{N}}$ is a solution of the noiseless operator equation $y^{0}=Kx$.
5 Practical Constraints and Training
5.1 Weakly Enforced Spectral Constraints
For convergence analysis, one key requirement is that the denoiser should take the form of a proximal step. For a gradient step denoiser $D_{\sigma}=\operatorname{Id}-\nabla g_{\sigma}$ to be a proximal operator $D_{\sigma}=\operatorname{prox}_{\phi_{\sigma}}$ of some weakly convex function $\phi_{\sigma}$, a sufficient condition is that $\nabla g_{\sigma}$ is $L_{\sigma}$-Lipschitz for some $L_{\sigma}<1$ gribonval2020characterization ; hurault2022proximal . Here, we use the subscript $\sigma$ to denote specifically that the denoiser is trained to remove Gaussian noise of standard deviation $\sigma$. As enforcing the Lipschitz condition through architectural choices or otherwise is difficult, a standard approach in practice is to penalize the network in the loss function if the Lipschitz constant is too large. Noting the equivalence of the Lipschitz constant and the spectral norm of $\nabla^{2}g_{\sigma}=J(\operatorname{Id}-D_{\sigma})$, this consists of adding a spectral regularization term of the form
$\mathbb{E}_{x\sim p,\xi\sim\mathcal{N}(0,\sigma^{2})}\max(\|J(\operatorname{Id}-D_{\sigma})(x+\xi)\|,1-\varepsilon).$
Here, $\varepsilon\in(0,1)$ is a tuneable hyperparameter to control how strongly the spectral constraint should be enforced. This penalizes the spectral norm of $\nabla^{2}g_{\sigma}$, typically approximated using a power iteration. This method can be extended also to learn monotone operators pesquet2021learning ; belkouchi2025learning , while other methods of softly enforcing the Lipschitz constant include (approximate) layer-wise projections onto the Stiefel manifold of orthogonal matrices anil2019sorting ; cisse2017parseval . However, as the Lipschitz constant is not strictly enforced to be less than one, the algorithms suffer from occasional divergence.
5.2 Backtracking for Lipschitz Control
As mentioned in Theorem˜4.6, the gradient-step paradigm for PnP instead replaces the gradient step in a splitting with a denoiser, and applies the proximal operator on the fidelity term gs_denoiser_hurault_2021 . In this case, the theoretically convergent sequence $F(x_{k})$ requires the computation of $F=f+\lambda g$. This is computable since $f$ is a known fidelity term, and $g$ takes the special form $g(x)=\frac{1}{2}\|x-N(x)\|$ for a neural network $N(x)$. In this case, since the step size in the splitting is allowed to be variable, it remains to find an upper bound on the Lipschitz constant of $\nabla g_{\sigma}$, such that $D_{\lambda,\sigma}=\operatorname{Id}-\lambda\nabla g_{\sigma}$ is a descent step.